assume L is between K & M

K-------------L--------------M
if L is the midpoint of KM(line over top) and KM = 5x-6 AND LM = 2x +4 what is the value of X?

SO:
K-------------L-------------M
5x-6 2x+4


Answer by sofiyac(983)

L is the midpoint so KL=LM
so 5x-6=2(2x+4) solve for x
5x-6=2x+8 subtract 2x from each side
3x-6=8 add 6 to each side
3x=14 divide each side by 3
x=14/3